OpenStudy (anonymous):

To solve for y in the equation 2x + y = 5, subtract 2 from both sides of the equation.

OpenStudy (anonymous):

True or false

OpenStudy (phi):

if we subtract 2 from both side: 2x+y-2= 5-2 or 2x+y-2= 3 does that look helpful? what about 2x+y-2x= 5-2x y+2x-2x= -2x+5 y= -2x+5